Solve the problem. You must write two equations with two variables and show steps to solve.
Anne and Nancy use a metal alloy that is 17% copper to make jewelry. How many ounces of an alloy that is 11% copper must be mixed with an alloy that is 22% copper to form 55 ounces of the desired alloy?

Answer by ikleyn(52776)

.
    x +     y =      55,   (1)
0.11x + 0.22y = 0.17*55    (2)


Can you guess what is x and what is y without my explanations?


To solve the system, express x = 55 - y from (1). Then substitute it into (2).
You will get a single equation for one unknown y

0.11*(55-y) + 0.22y = 9.35.

Simplify and solve for y.

Then calculate x from either of equations (1) or (2).

There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".